OpenStudy (anonymous):

Hi can someone please help me with the following two questions thanks. It is about calculus. Maybe if you show me the steps for the first one i can attempt the second.

OpenStudy (anonymous):

both require the chain rule

OpenStudy (anonymous):

the derivative of secant is secant times tangent and the derivative of \(\sqrt[3]{x}\) is \(\frac{1}{3\sqrt[3]{x^2}}\)

OpenStudy (anonymous):

your answer is therefore \[\sec(\sqrt[3]{x})\tan(\sqrt[3]{x})\times \frac{1}{3\sqrt[3]{x^2}}\]

OpenStudy (anonymous):

isn't the derivative of \[\sqrt[3]{x}\] is 1/3x^(2/3) thanks

OpenStudy (anonymous):

yes that is what i wrote too most times exponential notation is handy but if you want to evaluate anything you really need the radical notation

OpenStudy (anonymous):

oh ok thanks hahah

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i will attempt the second one now

OpenStudy (anonymous):

k you need the chain rule three twice for that one recall \(4\cos^3(\pi x)\) really means \(4\left(\cos(\pi x)\right)^3\)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes, and the final answer is \(-12~\pi~ sin(\pi x)~cos^2(\pi x)\)

OpenStudy (anonymous):

because \(\dfrac{d}{dx}x =1\)

OpenStudy (anonymous):

is their a simpilar was to do it? or a different way? thanks

OpenStudy (anonymous):

ok thanks but maybe u didn't understand me .. i just want to make sure i didn't skip steps or unknowingly used wrong connotations. even if there is a better way to re-arrange the method i used properly. Since i have to show full work. thanks